OpenStudy (1645323):

zero for this equation :y=x^2-6x-4

OpenStudy (1645323):

@yellowlegoguy99

OpenStudy (yellowlegoguy99):

Why me? :P

OpenStudy (1645323):

i don't know i just typed in a random username

OpenStudy (1645323):

you are the chosen one

OpenStudy (yellowlegoguy99):

:O

OpenStudy (1645323):

whats the zero?

OpenStudy (yellowlegoguy99):

Do you have any choices I can pick?

OpenStudy (1645323):

no

OpenStudy (yellowlegoguy99):

I'm only good with stuff that I can pick through, cause I understand it more, sorry

OpenStudy (1645323):

ok

OpenStudy (anonymous):

The zeroes are 6.61, and -0.61

OpenStudy (1645323):

@AnswerMyQuestions

OpenStudy (anonymous):

@1645323 I told you that the zeroes are 6.61 and -0.61

OpenStudy (1645323):

no they dont work

OpenStudy (anonymous):

x^2-6x-4 make a perfect square (x^2-2*x*3+3^2)-3^2-4=0 (x-3)^2-9-4=0 x-3)^2=13 x-3=sqrt13 x=3+sqrt13

OpenStudy (anonymous):

sqrt13= + - 3.605

OpenStudy (anonymous):

x=3+3.605=6.605 and x=3-3.605=-0.605

OpenStudy (anonymous):

@1645323

OpenStudy (1645323):

dude htey dont work i get -.003975 @g17

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

Those are the approximate answers.

OpenStudy (anonymous):

I rounded them.

OpenStudy (1645323):

after i put what he got into the equation that is my final answer

OpenStudy (anonymous):

x = -0.60555128 or 6.6055513

OpenStudy (anonymous):

The zeroes are REPEATING decimals.

OpenStudy (anonymous):

-0.003975 is wrong loll @1645323

OpenStudy (1645323):

what the crap. just tell me the rightr answer

OpenStudy (anonymous):

I did.

OpenStudy (anonymous):

There are 2 right answers: x = -0.60555128 or 6.6055513

OpenStudy (1645323):

ok